OpenStudy (anonymous):

What is the simplified form of y to the fifth power over x all over y-squared over x-squared ?

OpenStudy (anonymous):

\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

So would it be y^7/x^3

OpenStudy (anonymous):

\[\frac{ \frac{ y ^{5} }{ x ^{2} } }{ \frac{ y ^{2}} { x ^{2} } }\] is this what you started with?

OpenStudy (anonymous):

oops, x instead of x squared under y to the fifth

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, then look \[\frac{ \frac{ y ^{5} }{ x } }{ \frac{ y ^{2} }{ x ^{2} } }=\frac{ y ^{5} }{ x } \times \frac{ x ^{2} }{ y ^{2} }=?\]

OpenStudy (anonymous):

So you multiply across right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

you can cancel from top to bottom since it's all multiplication before multiplying across, if you want

OpenStudy (anonymous):

ohh so its xy^3

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

you're welcome!